# Spencer Unger, Narmada Varadarajan, Felix Weilacher

### v1

@@ Abstract.
We show that a circle and square of the same area in $\mathbb{R}^{2}$ are equidecomposable by translations using $\mathbf{\Delta}^{0}_{2}$ pieces. That is, pieces which are simultaneously $F_{\sigma}$ and $G_{\delta}$ sets. This improves a result of Máthé-Noel-Pikhurko and is the best possible complexity in terms of the Borel hierarchy. More generally we show that bounded sets $A,B\subseteq\mathbb{R}^{n}$ with small enough boundaries and the same nonzero Lebesgue measure are equidecomposable with pieces that are countable unions of finite Boolean combinations of translates of $A,B$, and open sets.
The improvement comes from constructions of low complexity toasts and related objects which should be independently useful within Borel combinatorics.
@@ 1.1. Circle squaring
Laczkovich [11] famously showed in 1990—answering a long-standing question of Tarski[17]—that a circle and square in $\mathbb{R}^{2}$ of the same area are equidecomposable by translations. That is, there is an $n\in\mathbb{N}$, vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{2}$, and a partition $A_{1},\ldots,A_{n}$ of the circle, so that $A_{1}+x_{1},\ldots,A_{n}+x_{n}$ is a partition of the square. Laczkovich’s result has been improved upon steadily over the last decade by a series of results showing that the pieces in this equidecomposition, originally chosen in a non-constructive way, can have stronger regularity properties. Grabowski, Máthé, and Pikhurko [7] showed that the pieces can be chosen to be Lebesgue measurable and to have the property of Baire. Marks and the first author [15] showed that the pieces can even be Borel; that is, a constructive equidecomposition of the circle and square exists.
Once we have Borel pieces, the Borel Hierarchy becomes a natural way to measure more finely just how constructive an equidecomposition is. Recall that in any topological space $X$, $\mathbf{\Sigma}^{0}_{1}(X)$ and $\mathbf{\Pi}^{0}_{1}(X)$ denote the collections of open and closed subsets of $X$ respectively. Then, for countable ordinals $\alpha<\omega_{1}$ we inductively define $\mathbf{\Sigma}^{0}_{\alpha}(X)$ as the collection of countable unions of sets in $\mathbf{\Pi}^{0}_{\beta}(X)$ for $\beta<\alpha$ and $\mathbf{\Pi}^{0}_{\alpha}(X)$ as the collection of complements of sets in $\mathbf{\Sigma}^{0}_{\alpha}(X)$. Finally, $\mathbf{\Delta}^{0}_{\alpha}(X)$ is the collection of sets which are in both $\mathbf{\Sigma}^{0}_{\alpha}(X)$ and $\mathbf{\Pi}^{0}_{\alpha}(X)$. For example, $\mathbf{\Sigma}^{0}_{2}(X)$ and $\mathbf{\Pi}^{0}_{2}(X)$ denote exactly the classes of $F_{\sigma}$ and $G_{\delta}$ sets respectively, and $\mathbf{\Delta}^{0}_{2}(X)$ denotes the collection of sets that are simulatneously $F_{\sigma}$ and $G_{\delta}$.
@@
Let $A,B\subseteq\mathbb{R}^{2}$ be a closed circle and square of the same area. $A$ and $B$ are equidecomposable by translations using $\mathbf{\Delta}^{0}_{2}$ pieces.
@@
Let $A,B\subseteq\mathbb{R}^{2}$ be a closed circle and square of the same area. Are $A$ and $B$ equidecomposable (by translations) using pieces that are Boolean combinations of open sets?
By a result of Dubins, Hirsch, and Karush [5] a circle and square are not scissors congruent, so the open sets involved in a positive answer to this question would still need to be somewhat complex.
@@
$\overline{\dim}_{\square}(A)$ $=\limsup_{\varepsilon\to 0}\frac{\log N(\varepsilon,S)}{\log(1/\varepsilon)}.$
Clearly, if $A\subseteq\mathbb{R}^{k}$ then $\overline{\dim}_{\square}(A)\leq k$. Even open sets in $\mathbb{R}^{k}$ can have boundaries with Minkowski dimension $k$, but, for instance, balls and cubes in $\mathbb{R}^{k}$ have boundaries with dimension k1k-1.
This condition turns out to be the main thing needed to generalize the circle-squaring results above: Laczkovich proved [12] that for any $k\in\mathbb{N}$, if $A,B\subseteq\mathbb{R}^{k}$ are Lebesgue measurable and bounded with $\overline{\dim}_{\square}(\partial A),\overline{\dim}_{\square}(\partial B)<k$ and $\lambda(A)=\lambda(B)>0$ ($\lambda$ denotes the Lebesgue measure), then $A$ and $B$ are equidecomposable by translations. The Grabowski–Máthé–Pikhurko result also applies in this case, as does the result of Marks and the first author if $A$ and $B$ are also Borel.
@@
Let $A,B\subseteq\mathbb{R}^{k}$ be bounded sets with $\lambda(A)=\lambda(B)>0$ and $\overline{\dim}_{\square}(\partial A),\overline{\dim}_{\square}(\partial B)<k$. Then, $A$ and $B$ are equidecomposable by translations with pieces that are countable unions of Boolean combinations of translates of $A$, $B$, and open sets. In particular, if $A,B\in\mathbf{\Delta}^{0}_{\alpha}$ for $\alpha\geq 2$, so are the pieces of the equidecomposition.
@@
    $\overline{\dim}_{\square}(A)=\overline{\dim}_{\square}(A^{\prime})$.
@@
    Let $\mathcal{A}\subseteq\mathcal{P}([0,1)^{k})$ Suppose $A^{\prime}$ and $B^{\prime}$ are equidecomposable by translations (in $\mathbb{T}^{k}$) using pieces that are images of sets in $\mathcal{A}$. Then $A$ and $B$ are equidecomposable by translations (in $\mathbb{R}^{k}$) using pieces that are Boolean combinations open sets and translates of sets in $\mathcal{A}$.
@@
Let $A,B\subseteq\mathbb{T}^{k}$ with $\lambda(A)=\lambda(B)>0$ and $\overline{\dim}_{\square}(\partial A),\overline{\dim}_{\square}(\partial B)<k$. Then, $A$ and $B$ are equidecomposable by translations using pieces that are countable unions of Boolean combinations of translates of AA,$B$, and open sets.
@@
Finally, we remark without proof that all the results of this paper can easily be made effective. For instance, in our equidecomposition of a circle and square in $\mathbb{R}^{2}$, we can supply for each piece $C$ a computable function $c$ from $\mathbb{N}^{2}$ to the set of open rectangles with rational coordinates such that $C=\bigcap_{n}\bigcup_{m}c(n,m)$. This is to say that Theorems 1.1, 1.4, and 1.6 can also be made optimal from the point of view of the so-called lightface Borel hierarchy.
@@ 1.2. Toasts of low complexity
This graph $G$ is the space in which all of the combinatorics of the proof from [15] take place. One of the main ingredients in their proof is the existence of a (Borel) witness to hyperfiniteness of $G$ with particularly nice geometric properties, called a toast. See Section 2.3 for a definition of toast and an overview of how it is used in the proof of Borel equidecomposition results. Toast was introduced and first constructed for Schreier graphs of Borel actions of $\mathbb{Z}^{d}$ by Gao, Jackson, Krohne, and Seward [6], and it has since become a central tool in descriptive combinatorics [^2] This may be a confusing statement given the date attached to [6]. Though this preprint was only recently released, its results and techniques have been circulated in the community for some time. .
In general a toast for a graph $G$ on $X$ is a sequence of sets $T_{0},T_{1},\ldots\subseteq X$ satisfying certain geometric properties. In the proof from [15], the complexity of the layers, $T_{n}$, has a large bearing on the complexity of the pieces of the final equidecomposition. When $X$ is Polish and the action on $X$ is continuous, it is easy to check that the toast constructed in [15] has layers which are Boolean combinations of $\mathbf{\Pi}^{0}_{3}$ sets. In particular they are $\mathbf{\Delta}^{0}_{4}$.
@@
The relationship between our toast construction and that of [6] is analogous to that between the construction in [16] and that in [15]: Though $\mathbb{T}^{k}$ is not zero dimensional, we are able to find modifications of parts of the [6] proof which save a level on the torus.
An important tool in our modification is the use of low-complexity witnesses to Borel asymptotic dimension, a tool introduced by Conley et al. [4] which has since become central in the study of Borel group actions and descriptive combinatorics. This leads us to the following generalization of Theorem 1.8.
@@
In particular, if the Bernoulli shift of $\Gamma$ admits clopen witnesses to asymptotic dimension, then any zero dimensional $\Gamma$-space admits toasts with $\mathbf{\Delta}^{0}_{2}$-layers, and any $\Gamma$-space admits toasts with $\mathbf{\Delta}^{0}_{3}$-layers.
Since toast has played a large role recently in many parts of descriptive combinatorics beyond just equidecomposition results, Theorem 1.9 may be of independent interest. This is helped by the fact that the hypothesis about clopen witnesses to asymptotic dimension has been been shown to hold for a wide variety of countable amenable groups [4, Theorem 10.7].
@@ 2.1. The Schreier graph
We fix the following parameters for the remainder of the paper. Let $\Gamma$ be a finitely generated group with symmetric generating set $S$ and $X$ a set on which $\Gamma$ acts. Often we will assume this action is free. $X$ will often carry a topology $\tau$, in which case we will usually assume the action is continuous. In this case $X$ is called a $\Gamma$-space. The Schreier graph of this action with respect to $S$ is the graph with vertex set $X$ and edges $(x,\gamma\cdot x)$ for $\gamma\in S$, $x\in X$, and will always be denoted $G$. The connected components of the graph $G$ are the orbits of the action $\Gamma\curvearrowright X$. We will use the following standard graph-theoretic notation:
@@ 2.2. Some complexity classes of sets
Let $\mathcal{A}\subseteq\mathcal{P}(X)$ be a collection of subsets of $X$. We define the following collections of sets, in increasing order of algorithmic complexity.
@@
As we have stated, in our main application $\mathbb{Z}^{d}\curvearrowright\mathbb{T}^{k}$ will be an action by independently randomly chosen translations $x_{1},\ldots,x_{d}\in\mathbb{T}^{k}$. With probability 1, each coordinate of each $x_{i}$ will be irrational, which makes it clear that any axis-parallel rectangle is pseudo-clopen. Thus $\mathbb{T}^{k}$ is almost surely pseudo-0-dimensional with respect to this action. In fact, it is pseudo-0-dimensional with respect to any countable group of translations by taking rectangles with an angle independent from the translations.
@@ A $q$-toast is a sequence $(T_{n})_{n\in\mathbb{N}}$ of subsets of $X$ (called layers) such that
A flow on $G$ is a function $\phi:E(G)\to\mathbb{R}$ satisfying $\phi(x,y)=-\phi(y,x)$ for all $(x,y)\in E(G)$. A flow is called integral if it takes values in $\mathbb{Z}$. For a flow $\phi$ on $G$, $\phi^{out}:X\to\mathbb{R}$ denotes the function $x\mapsto\sum_{y\in N(x)}\phi(x,y)$. For $S\subseteq X$ finite we also adopt the notation $\phi^{out}(S)=\sum_{x\in S}\phi^{out}(S)=\sum_{(x,y)\in\partial_{E}S}\phi(x,y)$. For $c:X\to\mathbb{R}$, $\phi$ is called a $c$-flow if $\phi^{out}=c$. A 0-flow is also called a circulation.
In the case $c=\mathbbm{1}_{A}-\mathbbm{1}_{B}$, a $c$-flow should be viewed as a sort of linearized equidecomposition from $A$ to $B$. This intuition is confirmed by Lemma 2.8 below.
For the rest of this section fix sets $A,B\subseteq\mathbb{T}^{k}$ satisfying the hypotheses of Theorem 1.6. Following [16, Section 2.4], Let
$\epsilon\in(0,k-\max\{\overline{\dim}_{\square}(\partial A),\overline{\dim}_{\square}(\partial B)\}),$
$d=\lfloor k/\epsilon\rfloor+1\in\mathbb{N},\textnormal{ and }\varepsilon=(d\epsilon-k)/k>0.$
[…]
@@
With probability one the following is true: For any bounded integral $(\mathbbm{1}_{A}-\mathbbm{1}_{B})$-flow $\phi$ on $G$, there is an equidecomposition between $A$ and $B$ whose pieces are local functions of $\phi,A,B$, and basic open rectangles.
@@
With probability 1, there is a $c>0$ and (bounded) flows $0=\phi_{0},\phi_{1},\phi_{2},\dots,$ on $G$ such that
@@
We now explain how toasts enter the picture. By (4), $\phi_{\infty}$ it is Borel if $A$ and $B$ are, but the integral flow theorem may not produce a Borel integral flow. Instead, Marks and Unger show that a toast with Borel layers can be used to “round” $\phi_{\infty}$ one layer at a time, producing successive Borel $\mathbbm{1}_{A}-\mathbbm{1}_{B}$-flows, say $\psi_{0},\psi_{1},\ldots$, with $\psi_{n}$ integral and equal to $\phi_{n+1}$ on the first $n$ layers. The pointwise limit of the $\psi_{n}$’s then exists and is a bounded integral Borel $\mathbbm{1}_{A}-\mathbbm{1}_{B}$-flow, proving the theorem.
@@
With probability one 1 the following is true: Let $T_{0},T_{1},\ldots$ be a 2-toast for the Schreier graph of $\mathbb{Z}^{d}\curvearrowright\mathbb{T}^{k}$. There is an equidecomposition from $A$ to $B$ with each piece in $\mathbf{\Delta}_{\Gamma}(\{A,B\}\cup\{T_{n}\mid n\in\mathbb{N}\}\cup\mathbf{\Sigma}^{0}_{1})$.

### v2

@@ Abstract.
We show that a disk and square of the same area in $\mathbb{R}^{2}$ are equidecomposable by translations using $\mathbf{\Delta}^{0}_{2}$ pieces. That is, pieces which are simultaneously $F_{\sigma}$ and $G_{\delta}$ sets. This improves a result of Máthé-Noel-Pikhurko and is the best possible complexity in terms of the Borel hierarchy. More generally we show that bounded sets $A,B\subseteq\mathbb{R}^{n}$ with small enough boundaries and the same nonzero Lebesgue measure are equidecomposable with pieces that are countable unions of finite Boolean combinations of translates of $A,B$, and open sets. The improvement comes from constructions of low complexity toasts and related objects which should be independently useful within Borel combinatorics.
@@ 1.1. Circle squaring
Laczkovich [11] famously showed in 1990—answering a long-standing question of Tarski[17]—that a (closed) disk and square in $\mathbb{R}^{2}$ of the same area are equidecomposable by translations. That is, there is an $n\in\mathbb{N}$, vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{2}$, and a partition $A_{1},\ldots,A_{n}$ of the disk, so that $A_{1}+x_{1},\ldots,A_{n}+x_{n}$ is a partition of the square. Laczkovich’s result has been improved upon steadily over the last decade by a series of results showing that the pieces in this equidecomposition, originally chosen in a non-constructive way, can have stronger regularity properties. Grabowski, Máthé, and Pikhurko [7] showed that the pieces can be chosen to be Lebesgue measurable and to have the property of Baire. Marks and the first author [15] showed that the pieces can even be Borel; that is, a constructive equidecomposition of the disk and square exists.
Once we have Borel pieces, the Borel Hierarchy becomes a natural way to measure more finely just how constructive an equidecomposition is. Recall that in any topological space $X$, $\mathbf{\Sigma}^{0}_{1}(X)$ and $\mathbf{\Pi}^{0}_{1}(X)$ denote the collections of open and closed subsets of $X$ respectively. Then, for countable ordinals $\alpha<\omega_{1}$ we inductively define $\mathbf{\Sigma}^{0}_{\alpha}(X)$ as the collection of countable unions of sets in $\mathbf{\Pi}^{0}_{\beta}(X)$ for $\beta<\alpha$ and $\mathbf{\Pi}^{0}_{\alpha}(X)$ as the collection of complements of sets in $\mathbf{\Sigma}^{0}_{\alpha}(X)$. Finally, $\mathbf{\Delta}^{0}_{\alpha}(X)$ is the collection of sets which are in both $\mathbf{\Sigma}^{0}_{\alpha}(X)$ and $\mathbf{\Pi}^{0}_{\alpha}(X)$. For example, $\mathbf{\Sigma}^{0}_{2}(X)$ and $\mathbf{\Pi}^{0}_{2}(X)$ denote exactly the classes of $F_{\sigma}$ and $G_{\delta}$ sets respectively, and $\mathbf{\Delta}^{0}_{2}(X)$ denotes the collection of sets that are simultaneously $F_{\sigma}$ and $G_{\delta}$.
@@
Let $A,B\subseteq\mathbb{R}^{2}$ be a closed disk and square of the same area. $A$ and $B$ are equidecomposable by translations using $\mathbf{\Delta}^{0}_{2}$ pieces.
@@
Let $A,B\subseteq\mathbb{R}^{2}$ be a closed disk and square of the same area. Are $A$ and $B$ equidecomposable (by translations) using pieces that are Boolean combinations of open sets?
By a result of Dubins, Hirsch, and Karush [5] a disk and square are never scissors congruent, so the open sets involved in a positive answer to this question would still need to be somewhat complex.
@@
$\overline{\dim}_{\textnormal{box}}(A)$ $=\limsup_{\varepsilon\to 0}\frac{\log N(\varepsilon,S)}{\log(1/\varepsilon)}.$
The Minkowski dimension is sometimes also called the box-counting dimension Clearly, if $A\subseteq\mathbb{R}^{k}$ then $\overline{\dim}_{\textnormal{box}}(A)\leq k$. Even open sets in $\mathbb{R}^{k}$ can have boundaries with Minkowski dimension $k$, but, for instance, balls and cubes in $\mathbb{R}^{k}$ have boundaries with dimension k1k-1.
This condition turns out to be the main thing needed to generalize the circle-squaring results above: Laczkovich proved [12] that for any $k\in\mathbb{N}$, if $A,B\subseteq\mathbb{R}^{k}$ are Lebesgue measurable and bounded with $\overline{\dim}_{\textnormal{box}}(\partial A),\overline{\dim}_{\textnormal{box}}(\partial B)<k$ and $\lambda(A)=\lambda(B)>0$ ($\lambda$ denotes the Lebesgue measure), then $A$ and $B$ are equidecomposable by translations. The Grabowski–Máthé–Pikhurko result also applies in this case, as does the result of Marks and the first author if $A$ and $B$ are also Borel.
@@
Let $A,B\subseteq\mathbb{R}^{k}$ be bounded sets with $\lambda(A)=\lambda(B)>0$ and $\overline{\dim}_{\textnormal{box}}(\partial A),\overline{\dim}_{\textnormal{box}}(\partial B)<k$. Then, $A$ and $B$ are equidecomposable by translations with pieces that are countable unions of Boolean combinations of translates of $A$, $B$, and open sets. In particular, if $A,B\in\mathbf{\Delta}^{0}_{\alpha}$ for $\alpha\geq 2$, so are the pieces of the equidecomposition.
@@
    $\overline{\dim}_{\textnormal{box}}(A)=\overline{\dim}_{\textnormal{box}}(A^{\prime})$.
@@
    Let $\mathcal{A}\subseteq\mathcal{P}([0,1/2]^{k})$ Suppose $A^{\prime}$ and $B^{\prime}$ are equidecomposable by translations (in $\mathbb{T}^{k}$) using pieces that are images of sets in $\mathcal{A}$. Then $A$ and $B$ are equidecomposable by translations (in $\mathbb{R}^{k}$) using pieces that are Boolean combinations open sets and translates of sets in $\mathcal{A}$.
@@
Let $A,B\subseteq\mathbb{T}^{k}$ with $\lambda(A)=\lambda(B)>0$ and $\overline{\dim}_{\textnormal{box}}(\partial A),\overline{\dim}_{\textnormal{box}}(\partial B)<k$. Then, $A$ and $B$ are equidecomposable by translations using pieces that are countable unions of Boolean combinations of translates of AA,$B$, and open sets.
@@
Finally, we remark without proof that all the results of this paper can easily be made effective. For instance, in our equidecomposition of a disk and square in $\mathbb{R}^{2}$, we can supply for each piece $C$ a computable function $c$ from $\mathbb{N}^{2}$ to the set of open rectangles with rational coordinates such that $C=\bigcap_{n}\bigcup_{m}c(n,m)$. This is to say that Theorems 1.1, 1.4, and 1.6 can also be made optimal from the point of view of the so-called lightface Borel hierarchy.
@@ 1.2. Toasts of low complexity
This graph $G$ is the space in which all of the combinatorics of the proof from [15] take place. One of the main ingredients in their proof is the existence of a (Borel) witness to hyperfiniteness of $G$ with special properties, called a toast. Toast was introduced and first constructed for Schreier graphs of Borel actions of $\mathbb{Z}^{d}$ by Gao, Jackson, Krohne, and Seward [6], and it has since become a central tool in descriptive combinatorics [^2] This may be a confusing statement given the date attached to [6]. Though this preprint was only recently released, its results and techniques have been circulated in the community for some time. .
In general a toast for a graph $G$ on $X$ is a sequence of sets $T_{0},T_{1},\ldots\subseteq X$ satisfying certain geometric properties (See Definition 2.6). In the proof from [15], the complexity of the layers, $T_{n}$, has a large bearing on the complexity of the pieces of the final equidecomposition. When $X$ is Polish and the action on $X$ is continuous, it is easy to check that the toast constructed in [15] has layers which are Boolean combinations of $\mathbf{\Pi}^{0}_{3}$ sets. In particular they are $\mathbf{\Delta}^{0}_{4}$.
@@
Like [16] did with the construction in [15], we will need to revisit the toast construction from [6] and introduce modifications which allow us to save a level on the torus. Our modification is based on the same topological properties of the torus as that in [16] (see Definition 2.3) but the way we utilize them is quite different. Our toast construction also works in a more general situation than free actions of $\mathbb{Z}^{d}$, as is explained in the following Theorem.
An important tool in our modification is the use of low-complexity witnesses to Borel asymptotic dimension, a tool introduced by Conley, Jackson, Marks, Seward, and Tucker-Drob [4] which has since become central in the study of Borel group actions and descriptive combinatorics. This leads us to the following generalization of Theorem 1.8.
@@
In particular, if the (free part of the) Bernoulli shift of $\Gamma$ admits clopen witnesses to asymptotic dimension, then any zero dimensional (free) $\Gamma$-space admits toasts with $\mathbf{\Delta}^{0}_{2}$-layers, and any $\Gamma$-space admits toasts with $\mathbf{\Delta}^{0}_{3}$-layers.
Since toast has played a large role recently in many parts of descriptive combinatorics beyond just equidecomposition results, Theorem 1.9 may be of independent interest. This is helped by the fact that the hypothesis about clopen witnesses to asymptotic dimension on the free part of the Bernoulli shift has been shown to hold for a wide variety of countable amenable groups, at least in the free case [4, Theorem 10.7].
@@ 2.1. The Schreier graph
We fix the following parameters for the remainder of the paper. Let $\Gamma$ be a finitely generated group with symmetric finite generating set $S$ and $X$ a set on which $\Gamma$ acts. Often we will assume this action is free. $X$ will often carry a topology $\tau$, in which case we will usually assume the action is continuous. In this case $X$ is called a $\Gamma$-space. The Schreier graph of this action with respect to $S$ is the graph with vertex set $X$ and edges $(x,\gamma\cdot x)$ for $\gamma\in S$, $x\in X$, and will always be denoted $G$. The connected components of the graph $G$ are the orbits of the action $\Gamma\curvearrowright X$. We will use the following standard graph-theoretic notation:
@@ 2.2. Some complexity classes of sets
Let $\mathcal{A}\subseteq\mathcal{P}(X)$ be a collection of subsets of $X$. We define the following collections of sets.
@@
As we have stated, in our main application $\mathbb{Z}^{d}\curvearrowright\mathbb{T}^{k}$ will be an action by independently randomly chosen translations $x_{1},\ldots,x_{d}\in\mathbb{T}^{k}$. With probability 1, the coordinates of the $x_{i}$’s together with 1 will be linearly independent over $\mathbb{Q}$, which makes it clear that any axis-parallel rectangle is pseudo-clopen. Thus $\mathbb{T}^{k}$ is almost surely pseudo-0-dimensional with respect to this action. In fact, it is pseudo-0-dimensional with respect to any countable group of translations by taking rectangles with an angle independent from the translations.
@@ A $q$-toast is a sequence $(T_{n})_{n\in\mathbb{N}}$ of subsets of $X$ (called layers) such that
A flow on $G$ is a function $\phi:E(G)\to\mathbb{R}$ satisfying $\phi(x,y)=-\phi(y,x)$ for all $(x,y)\in E(G)$. A flow is called integral if it takes values in $\mathbb{Z}$. For a flow $\phi$ on $G$, $\phi^{out}:X\to\mathbb{R}$ denotes the function $x\mapsto\sum_{y\in N(x)}\phi(x,y)$. For $S\subseteq X$ finite we also adopt the notation $\phi^{out}(S)=\sum_{x\in S}\phi^{out}(S)=\sum_{(x,y)\in\partial_{E}S}\phi(x,y)$. For $f:X\to\mathbb{R}$, $\phi$ is called an $f$-flow in $G$ if $\phi^{out}=f$. A 0-flow is also called a circulation.
When $G$ is the Schreier graph of an action of translations on the torus and $f=\mathbbm{1}_{A}-\mathbbm{1}_{B}$, an $f$-flow in $G$ should be viewed as a sort of linearized equidecomposition from $A$ to $B$. This intuition is confirmed by Lemma 2.11 below. The existence of an $f$-flow in $G$ is a consequence of discrepancy estimates for the sets $A$ and $B$ in the action generating $G$.
Definition 2.8.
For a finite $F\subseteq\mathbb{T}^{k}$ and Lebesgue measurable $A\subseteq\mathbb{T}^{k}$, let $D(A,F)=|\lambda(A)-\frac{|A\cap F|}{|F|}|$. We call $D(A,F)$ the discrepancy of $A$ with respect to $F$.
Definition 2.9.
Let $\mathbb{Z}^{d}\curvearrowright\mathbb{T}^{k}$ be an action by translations $x_{1},\dots,x_{d}$. For $x\in\mathbb{T}^{k}$ and $N\in\mathbb{N}$, let $R_{N}(x)=\{x+n_{1}x_{1}+\dots+n_{d}x_{d}\mid 0\leq n_{i}<N$ for all $i\leq n\}$.
Laczkovich ([12] Theorem 1) showed the following:
Theorem 2.10.
Let $x_{1},\dots x_{d}\in\mathbb{T}^{k}$ be such that the associated action $\mathbb{Z}^{d}\curvearrowright\mathbb{T}^{k}$ is free. If $A,B\subseteq\mathbb{T}^{k}$ have the same positive Lebesgue measure and there are constants $K$ and $\varepsilon$ such that for all $x\in\mathbb{T}^{k}$ and $N>0$, $D(A,R_{N}(x)),D(B,R_{N}(x))\leq KN^{-1-\varepsilon}$, then $A$ and $B$ are equidecomposible by translations.
Laczkovich also showed that for $A\subseteq\mathbb{T}^{k}$ with $\overline{\dim}_{\textnormal{box}}(\partial A)<k$, for all sufficiently large $d$, almost every choice of $x_{1},\dots x_{d}$ satisfies the hypothesis of the theorem. Subsequent results use the hypothesis of Theorem 2.10 to show that the pieces of the equidecomposition can be taken to be Lebesgue (or Baire) measurable [7] or Borel [15]. The proof in [16] uses the same discrepancy estimates with a random choice of translations, but interestingly the additional property of small boundary of the pieces uses a positive measure subset of choices of translations, see [16] Lemma 7.5. The paper [14] provides an alternative way to produce actions with the same discrepancy estimates. As an application, they produce such actions with algebraic irrational coordinates.
@@
Let $A,B\subseteq\mathbb{T}^{k}$ and $x_{1},\dots x_{d}\in\mathbb{T}^{k}$ satisfy the hypothesis of Theorem 2.10 and let $G$ be the Schreier graph of the associated action. For any bounded integral $(\mathbbm{1}_{A}-\mathbbm{1}_{B})$-flow $\phi$ on $G$, there is an equidecomposition between $A$ and $B$ whose pieces are local functions of $\phi,A,B$, and basic open rectangles.
@@
Let $A,B\subseteq\mathbb{T}^{k}$ and $x_{1},\dots x_{d}\in\mathbb{T}^{k}$ satisfy the hypothesis of Theorem 2.10 and let $G$ be the Schreier graph of the associated action. There are $\varepsilon,c>0$ and (bounded) flows $0=\phi_{0},\phi_{1},\phi_{2},\dots,$ on $G$ such that
@@
We now explain how toasts enter the picture. By (4), $\phi_{\infty}$ is Borel if $A$ and $B$ are, but the integral flow theorem may not produce a Borel integral flow. Instead, Marks and Unger show that a toast with Borel layers can be used to “round” $\phi_{\infty}$ one layer at a time, producing successive Borel $\mathbbm{1}_{A}-\mathbbm{1}_{B}$-flows, say $\psi_{0},\psi_{1},\ldots$, with $\psi_{n}$ integral and equal to $\psi_{n+1}$ on the first $n$ layers. The pointwise limit of the $\psi_{n}$’s then exists and is a bounded integral Borel $\mathbbm{1}_{A}-\mathbbm{1}_{B}$-flow, proving the theorem.
@@
Let $A,B\subseteq\mathbb{T}^{k}$ and $x_{1},\dots x_{d}\in\mathbb{T}^{k}$ satisfy the hypothesis of Theorem 2.10, let $G$ be the Schreier graph of the associated action and let $T_{0},T_{1},\ldots$ be a 2-toast for $G$. The sets $A$ and $B$ are equidecomposible with pieces in $\mathbf{\Delta}_{\Gamma}(\{A,B\}\cup\{T_{n}\mid n\in\mathbb{N}\}\cup\mathbf{\Sigma}^{0}_{1})$.